please help me solve -1/2 divided by 3/4
Answer by algebrahouse.com(1659)

-1/2 ÷ 3/4
= -1/2 x 4/3 {when dividing with fractions, multiply by the reciprocal}
= -4/6 {multiplied, horizontally, through numerators and denominators}
= -2/3 {reduced}

For more help from me, visit:
www.algebrahouse.com